OpenStudy (anonymous):

Ok, well -x^2 + 60x + 900 i can't really transform this into a vertex form some help please this is as far as i could go -(x^2-60x-900)

myininaya (myininaya):

\[-(x^2-60x)+900\] \[-(x^2-60x+(\frac{60}{2})^2)+900+(\frac{60}{2})^2\] \[-(x-30)^2+900+30^2\] \[-(x-30)^2+1800\]

OpenStudy (anonymous):

yeah half of -60 o=is -30 so write as \[y=-(x-30)^2+k\]

OpenStudy (anonymous):

then to find k you can do all the work my master myininaya did, or you can make the brilliant observation that if x = 30 you will get \[y=-(30-30)^2+k=k\] so just put in 30 for x in the original equation and see what you get

OpenStudy (anonymous):

o well actually at first it was x(60-x) = m than 60x-x^2=m than i got -x^2-60x+900-900 = m than - (x^2-60x-900) -900 = m sorry i didn't include that! :(

OpenStudy (anonymous):

in other words compute \[-30^2+60\times -30 +900\]

OpenStudy (anonymous):

what is m?

OpenStudy (anonymous):

m is maximum

OpenStudy (anonymous):

are you solving a word problem of some sort?

OpenStudy (anonymous):

so are you trying to find the maximum of \[x(60-x)\]?

OpenStudy (anonymous):

the problem is a word problem. The sum of two numbers is 60. Find the numbers if their product is a maximum X+Y=60 XY=M

OpenStudy (anonymous):

because if you are, by symmetry you should see that the max is when x = 30

OpenStudy (anonymous):

right

OpenStudy (anonymous):

of course. if you want a rectangle with largest area make a square

OpenStudy (anonymous):

\[M=x(60-x)=60x-x^2\] a parabola facing down. the vertex is at \[-\frac{b}{2a}=-\frac{60}{2}=30\]

myininaya (myininaya):

X+Y=60=>Y=60-X so M=XY=X(60-X)=60X-X^2 so we have M=-X^2+60X M=-(X^2-60X) \[M=-(X^2-60X+(\frac{60}{2})^2)+(\frac{60}{2})^2\] \[M=-(X-30)^2+900\]

OpenStudy (anonymous):

and if you replace x by 30 you get \[30\times 30=900\]

OpenStudy (anonymous):

i've said it before and i'll say it again: if i want the vertex of some quadratic i will compute \[-\frac{b}{2a}\] and then i have to complete no square or try to reconcile what i have with what i want. it completes the square for me

myininaya (myininaya):

lol

OpenStudy (anonymous):

really!

OpenStudy (anonymous):

automatic complete square machine. patent pending

OpenStudy (anonymous):

whats going on down here? lool

OpenStudy (anonymous):

btw buck it occurs to me that you have already solved the problem. when you asked "write it in vertex form" what you really meant was you had already figured out that the maximum area was 900 when x = 30

myininaya (myininaya):

\[P(x)=ax^2+bx+c=a(x^2+\frac{b}{a}x)+c\] \[=a(x^2+\frac{b}{a}x+(\frac{b}{2a})^2)+c-a*(\frac{b}{2a})^2\] \[=a(x+\frac{b}{2a})^2+c-\frac{b^2}{2a}\]

OpenStudy (anonymous):

the problem is a word problem. The sum of two numbers is 60. Find the numbers if their product is a maximum X+Y=60 XY=M this is the problem you wrote yes?

myininaya (myininaya):

oops i missed something

OpenStudy (anonymous):

however you got form there to \[-x^2+60+900\] meant you had already completed the problem. you got the 900 from somewhere and that is your answer

OpenStudy (anonymous):

:S WHATS HAPPENING?! :L

OpenStudy (anonymous):

im good, i solved it, thank you guys!

myininaya (myininaya):

vertex is \[(-\frac{b}{2a},c-\frac{b^2}{4a})\]

OpenStudy (anonymous):

your answer is: the the product is 900 and both numbers are 30

myininaya (myininaya):

\[P=a(x+\frac{b}{2a})^2+c-\frac{b^2}{4a}\]